### v2

@@ 4.3 Application to geometric graphs
In this section, we apply the bounds obtained in Theorem 3.11 to geometric graphs. Geometric graphs model the position in a $d$-dimensional space, so the result can be useful for controlling objects in the real world. This is a generalization of the result shown in [21].
@@
For such graphs, we can define the following parameters [21]:
@@
    $r=\underset{(u,v)\in\mathcal{E}}{\max}d_{\mathrm{E}}(u,v);$
@@
Proposition 4.6 (Theorem 4.1 in [21]).
Let $P\in\mathbb{R}^{n\times n}$ be a reversible consensus matrix with invariant measure $\bm{\pi}$, associated with a graph $\mathcal{G}=(V,\mathcal{E})$. Assume that all nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
Let $P\in\mathbb{R}^{n\times n}$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph of $P$. Assume that all the nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}(P)$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then
@@
and particularly, when $P^{*}P=PP^{*}$ holds then
@@
We omit the proof because it is analogous to Section 6 in [21] except the bounds of $J(P)$ should be evaluated by Theorem 3.11.
@@
This family of graphs is called a family of geometric graphs with bounded parameters, including grid graphs with boundaries where the edges are connected only to the nearest neighbors [21]. In this sense, this family can be understood as a generalization of the grid graph case.
Although Proposition 4.6 discuss the case of fixed geometric parameters, the similar result to Proposition 4.6 is also known for this family of geometric graphs with bounded parameters in [21]. Here, we state the generalized version of it as a corollary.
@@
Let $P_{n}\in\mathbb{R}^{n\times n}$ be a family of consensus matrices with invariant measure $\bm{\pi}_{n}$, and let $\mathcal{G}(P_{n})$ be the associated undirected graph of $P_{n}$. Assume that all nonzero entries of $P_{n}$ belong to the interval $p_{\min},p_{\max}]$ and all entries of $n\bm{\pi}$ belong to the interval $\bar{\pi}_{\min},\bar{\pi}_{\max}]$ and that the family of $\mathcal{G}(P_{n})$ is a family of geometric graphs with bounded parameters $(s_{n},r_{n},\gamma_{n},\rho_{n})$ and nodes lying in $\mathcal{Q}=[0,l_{n}]^{d}$ in which $\gamma_{n}<l_{n}/4$. Then
@@
and particularly, when $P^{*}P=PP^{*}$ holds then
@@
The proof can be done by combining the discussion of Theorem 4.7 and Section 4.3 in [21].
@@
For a family of geometric graphs with bounded parameters, $l_{n}$ must grow linearly to $n^{1/d}$ in $n\to\infty$ (Lemma 6.3 in [21]). This feature is used for setting of $l_{n}$ in the following numerical experiment.
@@
    • We change all nonzero entries in $P_{n}$ to random values chosen by a uniform distribution on $[b,1]$. Then we normalize each row of $P_{n}$ so that $P_{n}$ becomes a stochastic matrix. By this process, the elements of $P_{n}$ are guaranteed to be greater than or equal to $\frac{b}{b+\delta}$.
@@ Appendix A Calculation of $\gamma_{n}$ and $\rho_{n}$ in the numerical experiment
    We use Floyd-Warshall (see Section 23.2 in [8]) method to calculate the distance between all pairs of points. Then we calculate $\rho_{n}$ and judge whether $\rho_{n}<\rho$.

### v3

@@ 4.3 Application to geometric graphs
In this section, we apply the bounds obtained in Theorem 3.11 to geometric graphs. Geometric graphs model the position in a $d$-dimensional space, so the result can be useful for controlling objects in the real world. This is a generalization of the result shown in [mainp].
@@
For such graphs, we can define the following parameters [mainp]:
@@
    $r=\underset{(u,v)\in\mathcal{E}}{\max}d_{\mathrm{E}}(u,v),$
@@
Proposition 4.4 (Theorem 4.1 in [mainp]).
Let $P\in\mathbb{R}^{n\times n}$ be a reversible consensus matrix with invariant measure $\bm{\pi}$, associated with a graph $\mathcal{G}=(V,\mathcal{E})$. Assume that all nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then,
@@
Let $P\in\mathbb{R}^{n\times n}$ be a consensus matrix with invariant measure $\bm{\pi}$, and let $\mathcal{G}(P)$ be the associated undirected graph of $P$. Assume that all the nonzero entries of $P$ belong to the interval $p_{\min},p_{\max}]$and that $\mathcal{G}(P)$ is a geometric graph with parameters $(s,r,\gamma,\rho)$ and nodes lying in $\mathcal{Q}=[0,l]^{d}$ in which $\gamma<l/4$. Then,
@@
and particularly, if $P^{*}P=PP^{*}$ holds, then
@@
We omit the proof because it is analogous to Section 6 in [mainp] except the bounds of $J(P)$ should be evaluated by Theorem 3.11.
@@
This family of graphs is called a family of geometric graphs with bounded parameters, including grid graphs with boundaries where the edges are connected only to the nearest neighbors [mainp]. In this sense, this family can be understood as a generalization of the grid graph case.
Although Proposition 4.4 discuss the case of fixed geometric parameters, the similar result to Proposition 4.4 is also known for this family of geometric graphs with bounded parameters in [mainp]. Here, we state the generalized version of it as a corollary.
@@
Let $P_{n}\in\mathbb{R}^{n\times n}$ be a family of consensus matrices with invariant measure $\bm{\pi}_{n}$, and let $\mathcal{G}(P_{n})$ be the associated undirected graph of $P_{n}$. Assume that all nonzero entries of $P_{n}$ belong to the interval $p_{\min},p_{\max}]$ and all entries of $n\bm{\pi}$ belong to the interval $\bar{\pi}_{\min},\bar{\pi}_{\max}]$ and that the family of $\mathcal{G}(P_{n})$ is a family of geometric graphs with bounded parameters $(s_{n},r_{n},\gamma_{n},\rho_{n})$ and nodes lying in $\mathcal{Q}=[0,l_{n}]^{d}$ in which $\gamma_{n}<l_{n}/4$. Then,
@@
and particularly, if $P^{*}P=PP^{*}$ holds, then
@@
The proof can be done by combining the discussion of Theorem 4.5 and Section 4.3 in [mainp].
@@
For a family of geometric graphs with bounded parameters, $l_{n}$ must grow linearly to $n^{1/d}$ in $n\to\infty$ (Lemma 6.3 in [mainp]). This feature is used for setting of $l_{n}$ in the following numerical experiment.
@@
    • We change all nonzero entries in $P_{n}$ to random values chosen by a uniform distribution on $[b,1]$. Then, we normalize each row of $P_{n}$ so that $P_{n}$ becomes a stochastic matrix. By this process, the elements of $P_{n}$ are guaranteed to be greater than or equal to $\frac{b}{b+\delta}$.
@@ Appendix A Calculation of $\gamma_{n}$ and $\rho_{n}$ in the numerical experiment
    We use Floyd-Warshall (see Section 23.2 in [cormen2022introduction]) method to calculate the distance between all pairs of points. Then, we calculate $\rho_{n}$ and judge whether $\rho_{n}<\rho$.